In a local district 9800 are democrats, 11000 are republican and 1200 are independent. What is the probability that any of teh three partiteis would be selected at random?
Answer by ewatrrr(24785)

 
Hi,
In a local district 9800 are democrats, 11000 are republican and 1200 are independent.
22,000 in ALL.
P(democrat selected) = 9800/22000 = 98/220 = 49/110
P(republican selected) = 11000/22000 = 1/2
P(Independent selected) = 1200/22000 = 12/220 = 3/55